# Bochner-type Theorems for Distributional Category

### v1

@@ 1. Introduction
While Lusternik–Schnirelmann (LS) category was originally invented to provide an estimate for the number of critical points of a smooth function on a manifold, it soon became apparent that the concept connected in non-trivial ways with other topological invariants (see [CLOT03]). The more recent development of the idea of topological complexity and its relation to the motion planning problem (see [Far08]) spawned a re-evaluation of LS-category in that context, and this resulted in a new “categorical” notion called distributional category (see [DJ24] and [KW24], where the term analog category is used in the latter). LS-category and distributional category are most certainly different. For instance, the LS-category of any real projective space ${{\mathbb{R}}P^{n}}$ is $n$ while the distributional category is $1$ for all $n$. Yet, intriguingly, sometimes $\operatorname{{\mathsf{dcat}}}(X)=\operatorname{{\mathsf{cat}}}(X)$ (where we have used the obvious notations). In particular, from [KW24] (also see [Dr]), we know that for a torsion-free discrete group $\pi$ with finite cohomological (and homotopy) dimension $k$, we have $\operatorname{{\mathsf{dcat}}}(\pi)=\operatorname{{\mathsf{cat}}}(\pi)=k$. In [Ja25], the first author gave some conditions under which $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)$ holds when $M$ is a closed (not necessarily aspherical) manifold. In this paper, we would like to explore other situations where the above equality holds, or at least where natural topological or geometric conditions imply something interesting about $\operatorname{{\mathsf{dcat}}}$. We refer to these conditions as Bochner-type conditions for reasons that will become apparent below.
@@
The results of our paper suggest close relationships between distributional category and various classical invariants of interest, such as the first Betti number, the Gottlieb group, LS-category, macroscopic dimension, etc. Furthermore, $\operatorname{{\mathsf{dcat}}}$ affords us an opportunity to improve upon earlier results in the literature that connect these classical invariants (see, for example, [Op02a][Op02b][OS10], and [DDJ]). Our results are strongest in the setting of $c$-symplectic manifolds and non-negatively curved manifolds, where they also give us various new estimates and computations of $\operatorname{{\mathsf{dcat}}}$. In particular, unlike the results of [DJ24] and [Ja25], we obtain first examples of inessential manifolds $M$ (in the sense of [Gr83]) for which $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)$ holds due to various topological, group-theoretical, and geometric conditions rather than purely cohomological ones.
@@
Just as LS-category has found applications in many areas of topology, geometry, and critical point theory, we hope that the results of our paper will hint at similar interconnectedness of distributional category, and at least give some intuition about its behavior on spaces with extra structures.
Notations and conventions
In this paper, all topological spaces are assumed to be path-connected. We use the symbol “$=$” to denote homeomorphisms of spaces and isomorphisms of groups, and the symbol $\simeq$” to denote homotopy equivalences.
@@ 2.A. Topological basics
Recall that the $k$-th Betti number of $X$, denoted $b_{k}(X)$, is defined to be the rank of $H_{k}(X;{\mathbb{Z}})$. Equivalently, $b_{k}(X)=\dim_{{\mathbb{Q}}}(H_{k}(X;{\mathbb{Q}}))$. Let $R$ be a commutative ring with unity. Then the cup-length of the cohomology ring $H^{*}(X;R)$, denoted $c\ell_{R}(X)$, is defined to be the largest integer $k$ for which there exist $\alpha_{i}\in H^{s_{i}}(X;R)$, $1\leq i\leq k$, $s_{i}>0$, such that $\alpha_{1}\alpha_{2}\cdots\alpha_{k}\neq 0$. In this paper, we reserve the notation $\operatorname{{\mathsf{cup}}}(X)$ for the rational cup-length $c\ell_{{\mathbb{Q}}}(X)$. Let $\Gamma$ be a discrete group. Then its cohomological dimension, denoted $\operatorname{{\rm cd}}(\Gamma)$, is defined to be the largest integer $k$ such that $H^{k}(\Gamma,A)\neq 0$ for some ${\mathbb{Z}}\Gamma$-module $A$. For example, $\operatorname{{\rm cd}}({\mathbb{Z}}^{n})=n$ and if $\Gamma$ has torsion, then $\operatorname{{\rm cd}}(\Gamma)$ is infinite, [Br82]. The Eilenberg–Ganea theorem (see [Br82]) says that for $\operatorname{{\rm cd}}(\Gamma)=k\geq 3$, there is an Eilenberg–Mac Lane space $K(\Gamma,1)$ of dimension $k$, and this is the smallest such dimension.
@@ 2.B. LS-category
    If $Y\simeq X$, then $\operatorname{{\mathsf{cat}}}(Y)=\operatorname{{\mathsf{cat}}}(X).$
@@ 2.C. Distributional category
Recently, a probabilistic variant of LS-category was introduced in [DJ24] and [KW24]. This is the distributional category (or analog category in [KW24]). We refer to  [KW24] and [Ja25] for motivations behind this new invariant. Given a pointed metric space $(X,x_{0})$, the distributional category of $X$, denoted $\operatorname{{\mathsf{dcat}}}(X)$, is defined to be the smallest integer $n$ for which there exists a map $H:X\to{\mathcal{B}}_{n+1}(P_{0}(X))$ such that $H(x)(0)\in{\mathcal{B}}_{n+1}(P(x,x_{0}))$ for each $x\in X$. Here, $P_{0}(X)=\{\phi\in X^{[0,1]}\mid\phi(1)=x_{0}\}$ and $P(x,x_{0})=\{\phi\in P_{0}(X)\mid\phi(0)=x\}$ have the (metrizable) subspace topology coming from the compact-open topology of $X^{[0,1]}$, and ${\mathcal{B}}_{n+1}(Z)$ is the space of probability measures on $Z$ supported by at most $n+1$ points, equipped with the Lévy–Prokhorov metric, [Pr56]. For a proof of the following, we refer to [DJ24].
@@
Just like LS-category, $\operatorname{{\mathsf{dcat}}}(X)=0$ if and only if $X$ is contractible, and $\operatorname{{\mathsf{dcat}}}(S^{n})=1$. For each $n$, $\operatorname{{\mathsf{dcat}}}(T^{n})=\operatorname{{\mathsf{cat}}}(T^{n})=n$. If $X$ is a closed surface, then $\operatorname{{\mathsf{dcat}}}(X)=1$ if $X\simeq S^{2}$ or $X\simeq{\mathbb{R}}P^{2}$ and $\operatorname{{\mathsf{dcat}}}(X)=\operatorname{{\mathsf{cat}}}(X)=2$ otherwise.
In general, the distributional category and the LS-category of a space may differ. Indeed, $\operatorname{{\mathsf{dcat}}}({\mathbb{R}}P^{n})=1<n=\operatorname{{\mathsf{cat}}}({\mathbb{R}}P^{n})$ for each $n\geq 2$, and for each finite group $\Gamma$, we have that $\operatorname{{\mathsf{dcat}}}(\Gamma)<|\Gamma|<\infty=\operatorname{{\mathsf{cat}}}(\Gamma)$, see [DJ24][KW24]. The following is the main result of [KW24] (see also [Dr]).
@@
If $\Gamma$ is a torsion-free discrete group, then $\operatorname{{\mathsf{dcat}}}(\Gamma)=\operatorname{{\mathsf{cat}}}(\Gamma)=\operatorname{{\rm cd}}(\Gamma)$.
This result has an analog for non-aspherical spaces. A closed manifold $M$ with $\operatorname{{\mathsf{cat}}}(M)=\dim(M)$ is said to be essential [Gr83] (see also  [KR06]). In [Ja25], it was shown that an essential manifold $M$ has $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=\dim(M)$ if $\pi_{1}(M)$ is torsion-free and $M$ obeys an extra homological condition called the cap property. This result may be extended to (generalized) connected sums under certain “cap property” conditions (see [Ja25]).
@@ 2.D. $c$-symplectic manifolds
[…]
@@ 2.E. Gottlieb groups and Hurewicz rank
The Gottlieb group of a space $X$, denoted $G(X)$, is defined to be the subgroup of $\pi_{1}(X)$ consisting of the elements $\alpha\in\pi_{1}(X)$ that have associated maps $A:S^{1}\times X\to X$ with $A|_{S^{1}}\simeq\alpha$ and $A|_{X}\simeq 1_{X}$. This is equivalent to $G(X)$ being the image of the evaluation map $\text{ev}_{*}\colon\pi_{1}(X^{X},1_{X})\to\pi_{1}(X)$, where $\text{ev}(f)=f(x_{0})$ for some basepoint $x_{0}\in X$. Since $\text{ev}_{*}$ is the connecting map for the universal fibration $X\to\text{Baut}_{*}(X)\to\text{Baut}(X)$ (see [Got69]), we see that $G(X)\subset\mathcal{Z}\pi_{1}(X)$, where $\mathcal{Z}\pi_{1}(X)$ is the center of $\pi_{1}(X)$. In the above setting, if we replace $S^{1}$ by $S^{n}$ and $\pi_{1}$ by $\pi_{n}$, then we get the $n$-th Gottlieb group of $X$, denoted $G_{n}(X)$. The following are some basic properties of $G(X)$, see [Got69].
@@
Here is an example of the influence of the Gottlieb group on the structure of a space. Suppose $X$ is a topological space such that $H_{1}(X;{\mathbb{Z}})$ is finitely generated. Let $h:\pi_{1}(X)\to H_{1}(X;{\mathbb{Z}})$ be the Hurewicz homomorphism. Then the Hurewicz rank of $X$ is defined to be the number of ${\mathbb{Z}}$-summands of $H_{1}(X;{\mathbb{Z}})$ which are contained in $h(G(X))$. We refer to [Got89][Op02b] for a proof of the following.
@@ 3. Related results
In this section, we recall some results from the literature that we aim to improve and/or extend in this paper. The most basic result for us is the classical result of S. Bochner that a closed $n$-manifold $M$ with non-negative Ricci curvature has $b_{1}(M)\leq\dim(M)$ and, perhaps most importantly, if $b_{1}(M)=\dim(M)$, then $M$ must be a (flat) torus. The proof given by Bochner was analytic in nature, and a modern proof involving the Dirac operator may be found in [LM89]. In fact, a simpler proof follows from the powerful splitting theorem of Cheeger and Gromoll [CG71], which is stated as follows.
@@
If $M$ is a closed $n$-manifold with non-negative Ricci curvature, then there exists a finite covering space $M^{\prime}$ of $M$ and a diffeomorphism $M^{\prime}=T^{r}\times W$, where $W$ is a closed simply connected manifold.
In this paper, we will refer to the finite cover $M^{\prime}$ as a Cheeger–Gromoll splitting of $M$.
@@
If $\pi_{1}(M)$ is finite, then $M^{\prime}$ is just the universal cover of $M$. So, Theorem 3.1 is not interesting in this case. Hence, we will consider only those closed manifolds with non-negative Ricci curvature that have infinite fundamental groups.
Using the Cheeger–Gromoll splitting, Bochner’s theorem was improved by the second author in [Op02a] as follows.
@@
Let $M$ be a closed $n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}=T^{r}\times W$ is a Cheeger–Gromoll splitting of $M$, then
$b_{1}(M)\leq\operatorname{{\mathsf{cat}}}(M)-\operatorname{{\mathsf{cup}}}(W).$
Moreover, $b_{1}(M)=\operatorname{{\mathsf{cat}}}(M)$ if and only if $M$ is a torus. Furthermore, if $M$ is also a closed $c$-symplectic $2n$-manifold, then the Cheeger–Gromoll splitting takes the form $M^{\prime}=T^{2k}\times W$, where $W$ is a closed simply connected $c$-symplectic manifold, and
$b_{1}(M)\leq 2k\leq 2\operatorname{{\mathsf{cat}}}(M)-\dim(M).$
@@
The proof of Theorem 3.3 uses the fact that the homomorphism $p^{*}:H^{*}(M;{\mathbb{Q}})\to H^{*}(M^{\prime};{\mathbb{Q}})$ induced by the finite covering map $p:M^{\prime}\to M$ is a split monomorphism due to the corresponding transfer map in cohomology, see [Br82, Sections III.9 & III.10]. As a result, $b_{1}(M)\leq b_{1}(M^{\prime})$.
In the case when $M$ is a closed $c$-symplectic manifold, Theorem 3.3 uses the following lemma from [Op02a] which we shall use below.
@@
If $p:M^{\prime}\to M$ is a finite covering map and $M$ is a $c$-symplectic $2n$-manifold with a cohomology class $\omega\in H^{2}(M;{\mathbb{Q}})$, then $M^{\prime}$ is also $c$-symplectic with $\omega\in H^{2}(M^{\prime};{\mathbb{Q}})$. Moreover, if $M^{\prime}=T^{r}\times W$ for some closed simply connected manifold $W$, then
  1. (1)
    $r=2k$, and
  2. (2)
    $w=w_{1}\oplus w_{2}$ for cohomology classes $w_{1}\in H^{2}(T^{2k};{\mathbb{Q}})$ and $w_{2}\in H^{2}(W;{\mathbb{Q}})$ satisfying $w_{1}^{k}\neq 0$ and $w_{2}^{n-k}\neq 0$.
@@
The simplest example of a closed $c$-symplectic (actually symplectic) manifold having non-negative Ricci (actually sectional) curvature is $M=T^{2}\times S^{2}$, and here we have $b_{1}(M)=2<3=\operatorname{{\mathsf{cat}}}(M)$. Of course, this example generalizes to any product $T^{2k}\times N^{2n-2k}$, where $N$ is any simply connected symplectic $(2n-2k)$-manifold. For these examples, the second estimate of Theorem 3.3 is sharp — but the manifold is not a torus. Besides these product examples, there are the Calabi–Yau manifolds with infinite fundamental groups [Yau77] (although most physicists take their fundamental groups to be finite).
The paper [FH82], which formulated a minimal model approach to rational LS-category, essentially re-invigorated the subject. One of the main results of [FH82] relates the “size” of (all) the Gottlieb groups of a space with the LS-category of its rationalization (in the simply connected case) as follows: if $X$ is a simply connected topological space, then $\dim(G_{*}(X)\otimes_{{\mathbb{Z}}}{\mathbb{Q}})\leq\operatorname{{\mathsf{cat}}}(X_{{\mathbb{Q}}})$, where $X_{{\mathbb{Q}}}$ is the rationalization of $X$. For non-simply connected spaces, an integral analog of the above result for the first Gottlieb group was obtained in [Op02b] using Theorem 2.5.
@@
If $A$ is a finitely generated free abelian subgroup of $G(X)$, then
$\operatorname{{\mathsf{rank}}}(A)\leq\operatorname{{\mathsf{cat}}}(X).$
In particular, if $G(X)$ is finitely generated free abelian, then $\operatorname{{\mathsf{rank}}}(G(X))\leq\operatorname{{\mathsf{cat}}}(X)$. Moreover, if $X$ is a closed $n$-manifold, then equality holds if and only if $X\simeq T^{n}$.

### v2

@@ 1. Introduction
While Lusternik–Schnirelmann (LS) category was originally invented to provide an estimate for the number of critical points of a smooth function on a manifold, it soon became apparent that the concept connected in non-trivial ways with other topological invariants (see [CLOT03]). The more recent development of the idea of topological complexity and its relation to the motion planning problem (see [Far08]) spawned a re-evaluation of LS-category in that context, and this resulted in a new “categorical” notion called distributional category (see [DJ24] and [KW24], where the term analog category is used in the latter). LS-category and distributional category are most certainly different. For instance, the LS-category of any real projective space ${\mathbb{R}}P^{n}$} is $n$ while the distributional category is $1$ for all $n$. Yet, intriguingly, sometimes $\operatorname{{\mathsf{dcat}}}(X)=\operatorname{{\mathsf{cat}}}(X)$ (where we have used the obvious notations). In particular, from [KW24] (also see [Dra]), we know that for a torsion-free discrete group $\pi$ with finite cohomological (and homotopy) dimension $k$, we have $\operatorname{{\mathsf{dcat}}}(\pi)=\operatorname{{\mathsf{cat}}}(\pi)=k$. In [Jau25], the first author gave some conditions under which $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)$ holds when $M$ is a closed (not necessarily aspherical) manifold. In this paper, we would like to explore other situations where the above equality holds, or at least where natural topological or geometric conditions imply something interesting about $\operatorname{{\mathsf{dcat}}}$. We refer to these conditions as Bochner-type conditions for reasons that will become apparent below.
@@
The results of our paper suggest close relationships between distributional category and various classical invariants of interest, such as the first Betti number, the Gottlieb group, LS-category, macroscopic dimension, etc. Furthermore, $\operatorname{{\mathsf{dcat}}}$ affords us an opportunity to improve upon earlier results in the literature that connect these classical invariants (see, for example, [Opr02a][Opr02b][OS10], and [DCDJ]). Our results are strongest in the setting of $c$-symplectic manifolds and non-negatively curved manifolds, where they also give us various new estimates and computations of $\operatorname{{\mathsf{dcat}}}$. In particular, unlike the results of [DJ24] and [Jau25], we obtain first examples of inessential manifolds $M$ (in the sense of [Gro83]) for which $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)$ holds due to various topological, group-theoretical, and geometric conditions rather than purely cohomological ones.
@@
Just as LS-category has found applications in many areas of topology, geometry, and critical point theory, we hope that the results of our paper will hint at similar interconnectedness of distributional category, and at least give some intuition about its behavior on spaces with extra structures. We also mention several directions for further study that the reader may find interesting (see, for instance, Remark 5.6).
In this paper, all topological spaces are assumed to be path-connected.
@@ 2.A. Topological basics
Let $R$ be a commutative ring with unity. Then the cup-length of the cohomology ring $H^{*}(X;R)$, denoted $c\ell_{R}(X)$, is defined to be the largest integer $k$ for which there exist $\alpha_{i}\in H^{s_{i}}(X;R)$, $1\leq i\leq k$, $s_{i}>0$, such that $\alpha_{1}\alpha_{2}\cdots\alpha_{k}\neq 0$. In this paper, we reserve the notation $\operatorname{{\mathsf{cup}}}(X)$ for the rational cup-length $c\ell_{{\mathbb{Q}}}(X)$. Let $\Gamma$ be a discrete group. Then its cohomological dimension, denoted $\operatorname{{\mathsf{cd}}}(\Gamma)$, is defined to be the largest integer $k$ such that $H^{k}(\Gamma,A)\neq 0$ for some ${\mathbb{Z}}\Gamma$-module $A$. The Eilenberg–Ganea theorem and the Stallings–Swan theorem (see [Bro82]) together say that for $\operatorname{{\mathsf{cd}}}(\Gamma)=k\neq 2$, there is an Eilenberg–Mac Lane space $K(\Gamma,1)$ of dimension $k$, and this is the smallest such dimension.
@@ 2.B. LS-category
    If $Y$ is homotopy equivalent to $X$ (i.e., $Y\simeq X$), then $\operatorname{{\mathsf{cat}}}(Y)=\operatorname{{\mathsf{cat}}}(X).$
@@ 2.C. Distributional category
Recently, a probabilistic variant of LS-category was introduced in [DJ24] and [KW24]. This is the distributional category (or analog category in [KW24]). We refer to  [KW24] and [Jau25] for motivations behind this new invariant. Given a pointed metric space $(X,x_{0})$, the distributional category of $X$, denoted $\operatorname{{\mathsf{dcat}}}(X)$, is defined to be the smallest integer $n$ for which there exists a map $H:X\to{\mathcal{B}}_{n+1}(P_{0}(X))$ such that $H(x)(0)\in{\mathcal{B}}_{n+1}(P(x,x_{0}))$ for each $x\in X$. Here, $P_{0}(X)=\{\phi\in X^{[0,1]}\mid\phi(1)=x_{0}\}$ and $P(x,x_{0})=\{\phi\in P_{0}(X)\mid\phi(0)=x\}$ have the (metrizable) subspace topology coming from the compact-open topology of $X^{[0,1]}$, and ${\mathcal{B}}_{n+1}(Z)$ is the space of probability measures on $Z$ supported by at most $n+1$ points, equipped with the Lévy–Prokhorov metric, [Pro56]. We note that like $\operatorname{{\mathsf{cat}}}(X)$ (see [CLOT03, Lemma 1.25]), $\operatorname{{\mathsf{dcat}}}(X)$ is also independent of the choice of the basepoint $x_{0}\in X$, see also [DJ24, Definition 3.1].
For a proof of the following, we refer to [DJ24].
@@
For each $n$, $\operatorname{{\mathsf{dcat}}}(T^{n})=\operatorname{{\mathsf{cat}}}(T^{n})=n$. If $X$ is a closed surface, then $\operatorname{{\mathsf{dcat}}}(X)=1$ if $X\simeq S^{2}$ or $X\simeq{\mathbb{R}}P^{2}$, and $\operatorname{{\mathsf{dcat}}}(X)=\operatorname{{\mathsf{cat}}}(X)=2$ otherwise.
In general, the distributional category and the LS-category of a space may differ. Indeed, $\operatorname{{\mathsf{dcat}}}({\mathbb{R}}P^{n})=1<n=\operatorname{{\mathsf{cat}}}({\mathbb{R}}P^{n})$ for each $n\geq 2$, and for each finite group $\Gamma$, we have that $\operatorname{{\mathsf{dcat}}}(\Gamma)<|\Gamma|<\infty=\operatorname{{\mathsf{cat}}}(\Gamma)$, see [KW24] (and also [DJ24]). The following is the main result of [KW24] (see also [Dra]).
@@
If $\Gamma$ is a torsion-free discrete group, then $\operatorname{{\mathsf{dcat}}}(\Gamma)=\operatorname{{\mathsf{cat}}}(\Gamma)=\operatorname{{\mathsf{cd}}}(\Gamma)$.
In particular, $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=\dim(M)$ for closed aspherical manifolds (hence, nilmanifolds). This result has an analog for non-aspherical spaces. A closed manifold $M$ with $\operatorname{{\mathsf{cat}}}(M)=\dim(M)$ is said to be essential [Gro83] (see also  [KR06]). In [Jau25], it was shown that $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=\dim(M)$ for an essential manifold $M$ if $\pi_{1}(M)$ is torsion-free and $M$ obeys an extra homological condition called the cap property. This result may be extended to (generalized) connected sums under certain “cap property” conditions (see [Jau25]).
@@ 2.D. $c$-symplectic manifolds
A $2n$-dimensional manifold $M$ is called $c$-symplectic if there exists a cohomology class $\omega\in H^{2}(M;{\mathbb{Q}})$ such that $\omega^{n}\neq 0$. The appellation $c$-symplectic is used because these spaces “mimic” symplectic manifolds cohomologically, [LO95]. While each symplectic manifold is $c$-symplectic, the converse is not true. Indeed, ${\mathbb{C}}P^{2}\#{\mathbb{C}}P^{2}$ is $c$-symplectic but not symplectic. Some other examples of $c$-symplectic manifolds include the Calabi–Yau manifolds [Yau77] (which, of course, are Kähler) and symmetric products of closed orientable surfaces [Mac62] (see also [DCDJ]). If $M$ is a closed $2n$-symplectic manifold and $M$ is simply connected, then $\operatorname{{\mathsf{cup}}}(M)=\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=n$ holds due to Theorems 2.1 and 2.2. If $M$ is not simply connected, then both $\operatorname{{\mathsf{cat}}}$ and $\operatorname{{\mathsf{dcat}}}$ can attain any integer value between $n$ and $2n$. Indeed, for any $0\leq k\leq n$, one can start with a closed simply connected symplectic (2n2k(2n-2k)-manifold NN and create $M=N\times T^{2k}$ with $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=n+k$. For some special $c$-symplectic manifolds $M$, namely the symplectically aspherical manifolds, it was shown in [RO99] that $\operatorname{{\mathsf{cat}}}(M)=\dim(M)$. For a symplectically aspherical manifold $M$ satisfying the “cap property”, $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=\dim(M)$ was shown in [Jau25].
@@ 2.E. Gottlieb groups and Hurewicz rank
The Gottlieb group of a space $X$, denoted $G(X)$, is defined to be the subgroup of $\pi_{1}(X)$ consisting of the elements $\alpha\in\pi_{1}(X)$ that have associated maps $A:S^{1}\times X\to X$ with $A|_{S^{1}}\simeq\alpha$ and $A|_{X}\simeq 1_{X}$. This is equivalent to $G(X)$ being the image of the evaluation map $\text{ev}_{*}\colon\pi_{1}(X^{X},1_{X})\to\pi_{1}(X)$, where $\text{ev}(f)=f(x_{0})$ for some $x_{0}\in X$. Since $\text{ev}_{*}$ is the connecting map for the universal fibration $X\to\text{Baut}_{*}(X)\to\text{Baut}(X)$ (see [Got69]), we see that $G(X)\subset\mathcal{Z}\pi_{1}(X)$, where $\mathcal{Z}\pi_{1}(X)$ is the center of $\pi_{1}(X)$. In the above setting for $G(X)$, if we replace $S^{1}$ by $S^{n}$ and $\pi_{1}$ by $\pi_{n}$, then we get the $n$-th Gottlieb group of $X$, denoted $G_{n}(X)$. The following are some basic properties of $G(X)$, see [Got69].
@@
  5. (5)
    If $X\to E\to B$ is a fibration, then there is a map to the universal fibration (as above) which induces the commutative diagram
    ${\pi_{2}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial_{*}}$${\pi_{2}(\mathrm{Baut}(X))\cong\pi_{1}(X^{X},1_{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rm{ev}_{*}}$${\pi_{1}(X)}$
    that shows that $\operatorname{Im}(\partial_{*})\subset\operatorname{Im}(\mathrm{ev}_{*})\subset G(X)$.
Here is an example of the influence of the Gottlieb group on the structure of a space. Suppose $X$ is a topological space such that $H_{1}(X;{\mathbb{Z}})$ is finitely generated. Let $h:\pi_{1}(X)\to H_{1}(X;{\mathbb{Z}})$ be the Hurewicz homomorphism. Then the Hurewicz rank of $X$ is defined to be the number of $\mathbb{Z}$}-summands of $H_{1}(X;{\mathbb{Z}})$ which are contained in $h(G(X))$. We refer to [Got89][Opr02b] for a proof of the following.
@@ 3. Related results
In this section, we recall some results from the literature that we aim to improve and/or extend in this paper. The most basic result for us is the classical result of S. Bochner that a closed $n$-manifold $M$ with non-negative Ricci curvature has $b_{1}(M)\leq\dim(M)$ and, perhaps most importantly, if $b_{1}(M)=\dim(M)$, then $M$ must be (diffeomorphic to) a flat torus. The proof given by Bochner was analytic in nature, and a modern proof involving the Dirac operator may be found in [LM89]. In fact, a simpler proof follows from the powerful splitting theorem of Cheeger and Gromoll [CG71], which is stated as follows.
@@
If $M$ is a closed $n$-manifold with non-negative Ricci curvature, then there exists a finite covering space $M^{\prime}$ of $M$ and a diffeomorphism $M^{\prime}\cong T^{r}\times W$, where $W$ is a closed simply connected manifold.
We will refer to such a finite cover $M^{\prime}$ as a Cheeger–Gromoll splitting of $M$.
@@
If $\pi_{1}(M)$ is finite, then $M^{\prime}$ is the universal cover of $M$. So, Theorem 3.1 is not interesting in this case. Hence, we will consider only those closed manifolds with non-negative Ricci curvature that have infinite fundamental groups.
Using the Cheeger–Gromoll splitting, Bochner’s theorem was improved by the second author in [Opr02a] as follows.
@@
Let $M$ be a closed $n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}\cong T^{r}\times W$ is a Cheeger–Gromoll splitting of $M$, then $b_{1}(M)\leq\operatorname{{\mathsf{cat}}}(M)-\operatorname{{\mathsf{cup}}}(W)$. Moreover, $b_{1}(M)=\operatorname{{\mathsf{cat}}}(M)$ if and only if $M$ is a torus. Furthermore, if $M$ is also a closed $c$-symplectic $2n$-manifold, then the Cheeger–Gromoll splitting takes the form $M^{\prime}\cong T^{2k}\times W$, where $W$ is a closed simply connected $c$-symplectic manifold, and $b_{1}(M)\leq 2\operatorname{{\mathsf{cat}}}(M)-\dim(M)$.
@@
The proof of Theorem 3.3 uses the fact that the homomorphism $p^{*}:H^{*}(M;{\mathbb{Q}})\to H^{*}(M^{\prime};{\mathbb{Q}})$ induced by the finite covering map $p:M^{\prime}\to M$ is a split monomorphism due to the corresponding transfer map in cohomology, see [Bro82, Sections III.9 & III.10]. As a result, $b_{1}(M)\leq b_{1}(M^{\prime})$.
In the case when $M$ is a closed $c$-symplectic manifold, Theorem 3.3 uses the following lemma from [Opr02a], which we shall use below.
@@
If $p:M^{\prime}\to M$ is a finite covering map and $M$ is a $c$-symplectic $2n$-manifold with a cohomology class $\omega\in H^{2}(M;{\mathbb{Q}})$, then $M^{\prime}$ is also $c$-symplectic with $\omega\in H^{2}(M^{\prime};{\mathbb{Q}})$. Moreover, if $M^{\prime}\cong T^{r}\times W$ for some closed simply connected manifold $W$, then $r=2k$, and $w=w_{1}\oplus w_{2}$ for cohomology classes $w_{1}\in H^{2}(T^{2k};{\mathbb{Q}})$ and $w_{2}\in H^{2}(W;{\mathbb{Q}})$ satisfying $w_{1}^{k}\neq 0$ and $w_{2}^{n-k}\neq 0$.
@@
The simplest examples of closed $c$-symplectic (actually symplectic) manifolds having non-negative Ricci (actually sectional) curvature are the products $M^{2n}=T^{2k}\times N^{2n-2k}$ for closed simply connected symplectic (2n2k)(2n-2k)-manifolds NN. Here, we have $b_{1}(M)=2k<n+k=\operatorname{{\mathsf{cat}}}(M)$. For such a product manifold $M$, the second estimate of Theorem 3.3 is sharp, but $M$ is not a torus. Besides these examples, there are the Calabi–Yau manifolds with infinite fundamental groups [Yau77] (although most physicists take their fundamental groups to be finite).
The paper [FH82], which formulated a minimal model approach to rational LS-category, essentially re-invigorated the subject. One of the main results of [FH82] relates the “size” of (all) the Gottlieb groups of a space with the LS-category of its rationalization (in the simply connected case) as follows: if $X$ is a simply connected topological space, then $\dim(G_{*}(X)\otimes_{{\mathbb{Z}}}{\mathbb{Q}})\leq\operatorname{{\mathsf{cat}}}(X_{{\mathbb{Q}}})$, where $X_{{\mathbb{Q}}}$ is the rationalization of $X$. For non-simply connected spaces, an integral analog of the above result for the first Gottlieb group was obtained in [Opr02b] using Theorem 2.5.
@@
If $A$ is a finitely generated free abelian subgroup of $G(X)$, then $\operatorname{{\mathsf{rank}}}(A)\leq\operatorname{{\mathsf{cat}}}(X)$. In particular, if $G(X)$ is finitely generated free abelian, then $\operatorname{{\mathsf{rank}}}(G(X))\leq\operatorname{{\mathsf{cat}}}(X)$. Moreover, if $X$ is a closed $n$-manifold, then equality holds if and only if $X\simeq T^{n}$.